OpenStudy (anonymous):

Solve -7 < x + 2 < 4.

OpenStudy (anonymous):

-7<x+2 x>-9 x+2<4 x<2

OpenStudy (anonymous):

-9<x<2

OpenStudy (anonymous):

do you want number line ???

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

-7<x+2<4 =-7-2<x+2-2<4-2 =-9<x<2 Ans?